OpenStudy (anonymous):

The square of a number is equal to 10 less than 7 times that number. What are the two possible solutions? Which of the following equations is used in the process of solving this problem? x2 - 7x + 70 = 0 x2 - 7x + 10 = 0 x2 + 7x - 10 = 0

ganeshie8 (ganeshie8):

let the number be : \(x\) square of a number : \(x^2\) 7 times the number : \(7x\) 10 less above : \(7x-10\) so lets put the equation : \( x^2 = 7x-10\) now can you subtract or add... and make the RHS 0 ?

OpenStudy (anonymous):

x2=10-7x so making one side of the equation equals to 0; x2 + 7x -10 =0

OpenStudy (anonymous):

x^2 - 7x + 10?

ganeshie8 (ganeshie8):

correct :) now you can solve the quadratic equation to get the two possible solutions ! good luck

OpenStudy (anonymous):

thanks :)

ganeshie8 (ganeshie8):

welcome (: